OpenStudy (anonymous):

Ok, so I figured out a few problems, now I have more on the check list that I cannot understand. Like 100110011101 base two. WTH?

OpenStudy (anonymous):

Think about what our normal number system means 14792 in base ten represents \[14792 = 1*10^4 + 4*10^3 + 7*10^2 + 9*10^1 + 2*10^0\] So for base 2 100110011101 represents \[1*2^{11} + 0*2^{10} + 0*2^9 + 1*2^8 + 1*2^7 + 0*2^6 + 0*2^5 + 1*2^4 + 1*2^3 + 1*2^2 + 0*2^1 + 1*2^0\] \[= 2^{11} + 2^8 + 2^7 + 2^4 + 2^3 + 2^2 + 2^0\] \[= 2048 + 256 + 128 + 16 + 8 + 4 + 1\] \[= 2461\]

OpenStudy (anonymous):

Hrm.. part of the equation got cut off.

OpenStudy (anonymous):

100110011101 represents \[1*2^{11} + 0*2^{10} + 0*2^9 + 1*2^8 + 1*2^7 + 0*2^6 + 0*2^5 \] \[+ 1*2^4 + 1*2^3 + 1*2^2 + 0*2^1 + 1*2^0\] \[= 2^{11} + 2^8 + 2^7 + 2^4 + 2^3 + 2^2 + 2^0\] \[= 2048 + 256 + 128 + 16 + 8 + 4 + 1\] \[= 2461\]